# KnowVal: A Knowledge-Augmented and Value-Guided Autonomous Driving System

### v1

@@ Abstract
Visual–language reasoning, driving knowledge, and value alignment are essential for advanced autonomous driving systems. However, existing approaches largely rely on data-driven learning, making it difficult to capture the complex logic underlying decision-making through imitation or limited reinforcement rewards. To address this, we propose KnowVal, a new autonomous driving system that enables visual–language reasoning through the synergistic integration of open-world perception and knowledge retrieval. Specifically, we construct a comprehensive driving knowledge graph that encodes traffic laws, defensive driving principles, and ethical norms, complemented by an efficient LLM-based retrieval mechanism tailored for driving scenarios. Furthermore, we develop a human-preference dataset and train a Value Model to guide interpretable, value-aligned trajectory assessment. Experimental results show that our method substantially improves planning performance while remaining compatible with existing architectures. Notably, KnowVal achieves the lowest collision rate on nuScenes and state-of-the-art results on Bench2Drive.
@@ 1 Introduction
  • Extensive experiments show that KnowVal maintains strong compatibility with existing methods, and achieves the lowest collision rate on nuScenes and state-of-the-art performance on Bench2Drive.
@@ 2 Related Work
Atonomous Driving Systems. The prevailing paradigms for autonomous driving systems can be broadly categorized into three classes: the traditional modular pipeline, the end-to-end approach, and the Vision–Language–Action (VLA) model. Several models have been benchmarked on real-world datasets such as nuScenes [1]. ST-P3 [6] first integrates perception, prediction, and planning into an end-to-end framework. UniAD [7] follows a BEV-based feature extraction strategy similar to BEVFormer [15], with intricate feature propagation across task-specific decoders. VAD [12] adopts a fully vectorized scene representation for trajectory planning, while PARA-Drive [25] further investigates multi-task supervision and module interdependencies within end-to-end architectures. GenAD [32] utilizes BEV representations to obtain sparse detection and mapping features for planning, whereas SparseDrive [23] extracts such features directly from image inputs to mitigate cumulative errors and improve prediction accuracy. Recently, MomAD [22] introduces momentum-aware planning for enhanced temporal consistency, and BridgeAD [31] refines planning results by leveraging historical prediction and trajectory information. Finally, DiffusionDrive [16] employs a truncated diffusion model to generate diverse trajectory candidates, while HENet++ [28] jointly extracts foreground instance and dense panoramic features using attention mechanisms to improve prediction and planning.
@@ 4.1 Datasets and Metrics
We evaluate our approach on two widely used autonomous driving benchmarks: the open-loop real-world dataset nuScenes [1] and the closed-loop simulation benchmark Bench2Drive (B2D) [10]. nuScenes includes sensor data from six surround-view cameras, radar, and LiDAR across 1,000 driving scenes. Its planning metrics consist of the L2 error between predicted and human trajectories and the collision rate. Bench2Drive (B2D) provides 10,000 sampled driving segments for training and offers closed-loop evaluation environments with detailed simulator configurations. B2D employs two key metrics: Driving Score and Success Rate. The Driving Score is a composite measure that integrates Route Completion and Infraction Penalty, while the Success Rate reflects the percentage of successfully completed driving tasks.
4.2 Implementation Details
@@ 4.3 Main Results on Driving Benchmarks
Open-loop Evaluation on nuScenes. As reported in Table 1, KnowVal achieves the lowest Collision Rate among all existing methods on nuScenes. Although the L2 distance is slightly higher, this primarily indicates deviations from human driving trajectories rather than degraded planning quality. In practice, KnowVal tends to discover safer or more efficient driving strategies that may not be reflected in human demonstrations. These results also confirm that the KnowVal  can be seamlessly integrated into existing end-to-end architectures such as GenAD and HENet++, enhancing both safety and generalization.
Closed-loop Evaluation on Bench2Drive. As shown in Table 2, KnowVal achieves a further 3.35 improvement in Driving Score and a 1.76% gain in Success Rate over the VLA-based method SimLingo, establishing new state-of-the-art performance on Bench2Drive. These results demonstrate the effectiveness of our knowledge-grounded planning framework, enabling the autonomous driving system to make safer, more robust decisions in complex and dynamic driving environments.
@@ 4.5 Ablation
Ablation on Overall System. As shown in Table 3, integrating the knowledge retrieval and value model components, as well as incorporating open-world perception and retrieval-guided supplementary perception, consistently reduces the collision rate. Although the deviation from human driving trajectories increases slightly, this does not imply a decline in planning quality. We further analyze the effects of varying the numbers of knowledge entries and candidate trajectories. As shown in the middle and lower sections of the table, increasing these quantities beyond a certain threshold provides minimal additional benefit, indicating diminishing returns.
Table 3: Ablation studies on the KnowVal system. The baseline model is GenAD. K&V: using the proposed knowledge retrieval and value model. OW: incorporating open-ended perception and understanding of abstract concepts. RgP: using retrieval results to guide supplementary perception. $N_{K}$: the number of retrieved knowledge entries. $W_{K}$: applying weighting to the total score. $N_{T}$: the number of candidate trajectories.
@@
16 20 0.91 0.43
16 20 0.89 0.38
16 20 1.47 0.34
16 20 1.51 0.33
16 10 1.56 0.37
16 20 1.51 0.33
16 30 1.52 0.33
8 20 1.29 0.36
16 20 1.67 0.38
16 20 1.51 0.33
24 20 1.49 0.34
Table 4: Comparison on different Value Model architectures.
@@
BEV Image Total Score 10K 0.35 0.59
BEV Image Total Score 128K 0.22 0.44
Scene Feature Total Score 128K 0.13 0.27
Scene Feature Subscores 128K 0.11 0.20
@@ 5 Conclusion
A Qualitative Analysis of the Value Model
B Additional Implementation Details
B.1 Prompt used for scoring trajectories with MLLM
@@ 2. A "reasoning" object containing four keys:
B.2 Prompt for keyword extraction in the retrieval stage
@@ 5. When finished, output {completion_delimiter}
C About the Knowledge Graph
@@
D Sample of the Value Model Dataset

### v2

@@ Abstract
Visual–language reasoning, driving knowledge, and value alignment are essential for advanced autonomous driving systems. However, existing approaches largely rely on data-driven learning, making it difficult to capture the complex logic underlying decision-making through imitation or limited reinforcement rewards. To address this, we propose KnowVal, a new autonomous driving system that enables visual–language reasoning through the synergistic integration of open-world perception and knowledge retrieval. Specifically, we construct a comprehensive driving knowledge graph that encodes traffic laws, defensive driving principles, and ethical norms, complemented by an efficient LLM-based retrieval mechanism tailored for driving scenarios. Furthermore, we develop a human-preference dataset and train a Value Model to guide interpretable, value-aligned trajectory assessment. Experimental results show that our method substantially improves planning performance while remaining compatible with existing architectures. Notably, KnowVal achieves the lowest collision rate on nuScenes and state-of-the-art results on Bench2Drive and NVISIM.
@@ 1 Introduction
  • Extensive experiments show that KnowVal maintains strong compatibility with existing methods, and achieves the lowest collision rate on nuScenes and state-of-the-art performance on Bench2Drive and NVISIM.
@@ 2 Related Work
Atonomous Driving Systems. The prevailing paradigms for autonomous driving systems can be broadly categorized into three classes: the traditional modular pipeline, the end-to-end approach, and the Vision–Language–Action (VLA) model. Several models have been benchmarked on real-world datasets such as nuScenes [1] and NAVSIM [3]. ST-P3 [9] first integrates perception, prediction, and planning into an end-to-end framework. UniAD [10] follows a BEV-based feature extraction strategy similar to BEVFormer [18], with intricate feature propagation across task-specific decoders. VAD [15] adopts a fully vectorized scene representation for trajectory planning, while PARA-Drive [29] further investigates multi-task supervision and module interdependencies within end-to-end architectures. GenAD [36] utilizes BEV representations to obtain sparse detection and mapping features for planning, whereas SparseDrive [27] extracts such features directly from image inputs to mitigate cumulative errors and improve prediction accuracy. Recently, MomAD [26] introduces momentum-aware planning for enhanced temporal consistency, and BridgeAD [35] refines planning results by leveraging historical prediction and trajectory information. Finally, DiffusionDrive [20] employs a truncated diffusion model to generate diverse trajectory candidates, while HENet++ [32] jointly extracts foreground instance and dense panoramic features using attention mechanisms to improve prediction and planning.
@@ 4.1 Datasets and Metrics
We evaluate our approach on three widely used autonomous driving benchmarks: the real-world benchmark nuScenes [1] and NAVSIM [3], and the simulation benchmark Bench2Drive [13]. nuScenes [1] includes sensor data from six surround-view cameras, radar, and LiDAR across 1k driving scenes, about 35k samples. Its open-loop planning metrics consist of the L2 error between predicted and human trajectories and the collision rate. Bench2Drive [13] provides 10k sampled driving segments for training and offers closed-loop evaluation environments with detailed simulator configurations. Bench2Drive employs two key metrics: Driving Score (combining Route Completion and Infraction Penalty) and Success Rate. NAVSIM [3] introduces closed-loop metrics for open-loop evaluation, including the PDM score (PDMS) combining sub-metrics (NC, DAC, EP, TTC, C) over a 4-second non-reactive simulation. It includes 103k training and 12k testing samples.
@@ 4.2 Main Results on Driving Benchmarks
Table 3: Open-loop results with closed-loop metrics on NVISIM.
[…]
VADv2 [2] LC 97.2 89.1 91.6 100 76.0 80.9
Hydra-MDP [19] LC 98.3 96.0 94.6 100 78.7 86.5
DiffusionDrive [20] LC 98.2 96.2 94.7 100 82.2 88.1
KnowVal-DiffusionDrive LC 99.0 98.3 96.3 100 84.3 90.9 +2.8
UniAD [10] C 97.8 91.9 92.9 100 78.8 83.4
PARA-Drive [29] C 97.9 92.4 93.0 99.8 79.3 84.0
iPad [7] C 98.6 98.3 94.9 100 88.0 91.7
KnowVal-iPad C 99.2 98.7 95.4 100 88.1 92.9 +1.2
Open-loop Evaluation on nuScenes. As reported in Table 1, KnowVal achieves the lowest Collision Rate among all existing methods on nuScenes. Although the L2 distance is slightly higher, this primarily indicates deviations from human driving trajectories rather than degraded planning quality. In practice, KnowVal tends to discover safer or more efficient driving strategies that may not be reflected in human demonstrations.
Closed-loop Evaluation on Bench2Drive. As shown in Table 2, KnowVal achieves a further 3.35 improvement in Driving Score and a 1.76% gain in Success Rate over the VLA-based method SimLingo, establishing new state-of-the-art performance on Bench2Drive.
Open-loop results with closed-loop metrics on NVISIM. As shown in Table 3, to further demonstrate the compatibility of KnowVal with existing methods, we selected the advanced DiffusionDrive [20] and iPad [7] methods for experiments on NAVSIM. Both methods inherently support diverse candidate trajectory generation, thus eliminating the need for the modifications described in Section  3.3.1, and open-world perception, retrieval, and the value model can be directly integrated on them.
For more implementation details, please refer to Appendix A.
Experiment results above demonstrate the effectiveness of our knowledge-grounded planning framework, and show that KnowVal can be seamlessly integrated into existing end-to-end architectures, enhancing both safety and generalization.
@@ 4.4 Ablation
As shown in Table 4 and Table 5, integrating the knowledge retrieval and value model components, as well as incorporating open-world perception and retrieval-guided supplementary perception, consistently reduces the collision rate. Although the deviation from human driving trajectories increases slightly, this does not imply a decline in planning quality. We also analyze the effects of varying the numbers of knowledge entries and candidate trajectories in Appendix C. And we evaluate the robustness to noisy knowledge items in Appendix C.
Table 4: Ablation studies on nuScenes. The baseline model is GenAD. K&V: using the proposed knowledge retrieval and value model. OW: incorporating open-ended perception and understanding of abstract concepts. RgP: using retrieval results to guide supplementary perception.
@@
0.91 0.43
0.89 0.38
1.47 0.34
1.51 0.33
Table 5: Ablation studies on Bench2Drive. The baseline model is SimLingo. OW: incorporating open-ended perception and understanding of abstract concepts. RgP: using retrieval results to guide supplementary perception.
@@
A 85.07 67.27
B 84.91 67.03
C 85.92 67.69
D Heuristic Rules [27] 85.94 67.82
E 87.61 68.40
F 88.16 68.85
G 88.42 69.03
@@ 5 Conclusion
Supplementary Material
A Implementation Details
A.1 Training and Architecture Details
A.2 Prompt used for scoring trajectories with MLLM
@@ 2. A "reasoning" object containing four keys:
A.3 Prompt for keyword extraction in the retrieval stage
@@ 5. When finished, output {completion_delimiter}
B About the Knowledge Graph
@@
C Additional Ablation
Ablation on nuScenes. As shown in Table 6, integrating the knowledge retrieval and value model components, as well as incorporating open-world perception and retrieval-guided supplementary perception, consistently reduces the collision rate. Although the deviation from human driving trajectories increases slightly, this does not imply a decline in planning quality. We further analyze the effects of varying the numbers of knowledge entries and candidate trajectories. As shown in the middle and lower sections of the table, increasing these quantities beyond a certain threshold provides minimal additional benefit, indicating diminishing returns.
Ablation on Bench2Drive. As shown in Table 7, The experiments show each component’s effectiveness and robustness to noisy knowledge items.
Table 6: Ablation studies on nuScenes. The baseline model is GenAD. K&V: using the proposed knowledge retrieval and value model. OW: incorporating open-ended perception and understanding of abstract concepts. RgP: using retrieval results to guide supplementary perception. $N_{K}$: the number of retrieved knowledge entries. $W_{K}$: applying weighting to the total score. $N_{T}$: the number of candidate trajectories.
[…]
16 20 0.91 0.43
16 20 0.89 0.38
16 20 1.47 0.34
16 20 1.51 0.33
16 10 1.56 0.37
16 20 1.51 0.33
16 30 1.52 0.33
8 20 1.29 0.36
16 20 1.67 0.38
16 20 1.51 0.33
24 20 1.49 0.34
Table 7: Ablation studies on Bench2Drive. Exp. A is the SimLingo baseline. Exp. B uses the retrieved language tokens as Key-Value pairs for the planning transformer. Exp. C trains the value model using only the future-state tokens $S_{i}$ as input—excluding knowledge entries—while keeping the training setting consistent. Exp. D uses heuristic rules (SparseDrive; Sun et al., 2025) rather than a learned value model. Exp. H randomly inserts $N_{k}$ random knowledge items into the retrieved entries, while Exp. I uses $N_{k}$ randomly selected knowledge items without any retrieval. The experiments show each component’s effectiveness and robustness to noisy knowledge items.
ID Knowledge Values OW RgP DS SR
A 85.07 67.27
B 84.91 67.03
C 85.92 67.69
D Heuristic Rules 85.94 67.82
E 87.61 68.40
F 88.16 68.85
G 88.42 69.03
H Noise 87.07 67.99
I Random 86.18 67.74
J 87.61 68.40
Table 8: Comparison on different Value Model architectures.
[…]
BEV Image Total Score 10K 0.35 0.59
BEV Image Total Score 128K 0.22 0.44
Scene Feature Total Score 128K 0.13 0.27
Scene Feature Subscores 128K 0.11 0.20
D Analysis of Retrieval Validity
We conducted statistics on nuScenes val. 50.1% of the retrieved items are Key Items (which certainly fit in this scenario), 41.8% of them are General Items (fit in any scenario, e.g., speed limits), and 8.1% of them are Irrelevant Items (e.g., ‘give way to ambulance’, but there is no ambulance). Crucially, the fraction of Irrelevant Items did not exceed 2/16 in any single trajectory. The rows H to J in Table 7 also demonstrate robustness to noisy entries, as we train the value model to score 0 for irrelevant entries.
E Qualitative Analysis of the Value Model
F Sample of the Value Model Dataset